# Serrin’s overdetermined theorem within Lipschitz domains

### v3

@@ 1.2. Wulff potential and nontangential limits
We say that $Du$ is $\omega_{0}$-$L^{2}$-Dini-VMO up to the boundary for some $\omega_{0}>0$, if for any $x\in\Omega$ and $0<r<\frac{d(x)}{2}$,
[…]
satisfies
$\int_{0}^{1}r^{-1}{\sup_{\begin{subarray}{c}x\in\Omega,\\ {\mathop{\mathrm{\,dist\,}}}(x,\,\partial\Omega)=2r\end{subarray}}w^{2}(x,r)}\,dr\leq\omega_{0}$
and $w(x,\,r)\to 0$ as $x\to\partial\Omega$. One can generalize our approach for Theorem 1.1 to establish an anisotropic variant of Serrin’s overdetermined theorem for Lipschitz domains, with additional assumptions on $u$.
@@
Let $\Omega$ be a Lipschitz domain with a local Lipschitz constant $L>0$, $K$ be a (bounded) uniformly convex set of $C^{3}$ class with the principle curvatures of $\partial K$ bounded from above by $\kappa$ and from below by $\kappa^{-1}>0$, and $H$ be the associated Wulff potential. Assume that
$u\in W^{1,\,2}(\mathbb{R}^{n})\cap C^{2}_{\mathop{\mathrm{\,loc\,}}}(\Omega)\cap W^{2,\,2}(\Omega)$
satisfies the anisotropic overdetermined system (1.4) with $Du\in VMO(\Omega;\,\mathbb{R}^{n})$. Moreover, in a neighborhood of $\partial\Omega$ we assume that $D^{2}u\in L^{n}$, $|Du|\geq c_{0}>0$ and $Du$ is $\omega_{0}$-$L^{2}$-Dini-VMO up to the reduced boundary $\partial^{*}\Omega$ (outside a set of $\mathscr{H}^{n-1}$-measure zero). Then there exists a constant $\epsilon_{0}=\epsilon_{0}(n,\,\kappa,\,\|K\|_{C^{3}})>0$ satisfying that, whenever
$\max\{L,\,\omega_{0}\}\leq\epsilon_{0},$
one has
@@ 1.3. Alexandrov theorem and Maggi’s conjecture
As $\Omega_{m}$ is $C^{2}$, the Hopf boundary lemma implies
@@ 3. Preliminaries for the anisotropic case
$d\mu_{\mathcal{A}}(x)=\chi_{\Omega}(x)\sup\{d(z)|D\mathcal{A}(z)|^{2}\colon z\in B_{d(x)/4}(x)\}d\mathcal{L}(x)$
is a $C$-Carleson measure with respect to $\mathscr{H}^{n-1}|_{\partial\Omega}$, i.e., for any ball $B_{r}(y)$ centered at $y\in\partial\Omega$,
@@
$\mu_{\mathcal{A}}(B_{r}(y)\cap\Omega)\leq C\omega_{0}r^{n-1},$
where
$0<C=C(n,\,c_{0},\,L,\,\kappa,\,\|D^{2}V\|_{C^{\alpha}(\mathbb{S}^{n-1})},\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})}).$
@@
Let $y\in\partial^{*}\Omega$ and $x\in B_{r}(y)$. By differentiating the equation in (1.4) in the $e_{k}$-direction, $1\leq k\leq n$, and letting $v=\partial_{k}u$, we obtain that $v$ satisfies, in the sense of distribution,
@@
where $\kappa$ is the curvature bound of $\partial K$ in Theorem 1.2. Then by summing over all $k$’s, as $|Du|\neq 0$ in a neighborhood of $\partial\Omega,$ it follows from the Schauder estimates in divergence form that, for any $0<t\leq\frac{d(x)}{4}$ with $B_{t}(x)$ disjoint from $\Omega_{0}\subset\subset\Omega$,
@@
and the 1-1-homogeneity of $D^{3}V$ further gives that, for any $z\in B_{{d(x)}/4}(x)$,
@@
with uniformly bounded overlaps up to $C(n)$, covering $B_{r}(y)\cap\Omega$ so that
@@
$\leq$ $\ C(n,\,c_{0},\,\alpha,\,\|D^{2}V\|_{C^{\alpha}(\mathbb{S}^{n-1})},\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})})\sum_{i}\omega^{2}(x_{i},\,d(x_{i})/4)(d(x_{i}))^{n-1}$
$\leq$ $\ C(n,\,c_{0},\,L,\,\alpha,\,\|D^{2}V\|_{C^{\alpha}(\mathbb{S}^{n-1})},\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})})\omega_{0}r^{n-1},$
where in the last inequality we used the $\omega_{0}$-$L^{2}$-Dini-VMO of $Du$ at $y$, and that there are at most $C(n,\,L)2^{l(n-1)}r^{n-1}$-many balls in the collection with radius between $2^{-l}$ and $2^{-l+1}$ with $l\in\mathbb{N}$. Thus, we obtain the desired estimate in the case $y\in\partial^{*}\Omega.$
When $y\in\partial\Omega\setminus\partial^{*}\Omega$, as $\Omega$ is Lipschitz, we can find a point $\hat{y}\in\partial^{*}\Omega$ so that
$B_{r}(y)\cap\Omega\subset B_{2r}(\hat{y})\cap\Omega.$
This proves the lemma. ∎
@@
Recall that $Du\in VMO(\Omega;\,\mathbb{R}^{n})$. Then the assumption of Theorem 1.2 yields the existence of $\rho>0$ for which
$D^{2}u\in L^{n}(N_{\rho}(\partial\Omega)\cap\Omega)\quad\text{ and }\quad|Du|(z)\geq c_{0}>0\ \text{ for $z\in N_{\rho}(\partial\Omega)\cap\Omega$},$
@@
$N_{\rho}(\partial\Omega):=\{z\in\mathbb{R}^{n}\colon{\mathop{\mathrm{\,dist\,}}}(z,\,\partial\Omega)<\rho\}.$
@@
$\mathcal{A}(x)=D^{2}V(Du)\in VMO(\Omega),\quad D\mathcal{A}(z)=D^{3}V(Du(z))D^{2}u(z)\in L^{n}(N_{\rho}(\partial\Omega)\cap\Omega)$
to a global VMO vector field with $D\mathcal{A}(z)\in L^{n}(N_{\rho}(\partial\Omega))$, and still denote the extended one by $\mathcal{A}$.
@@
Then
${\rm div}({\mathcal{A}}(x)Dv)=\mathcal{A}D^{2}v+{\rm tr}(D\mathcal{A})Dv$
is uniformly elliptic, and $v\in W^{2,\,2}(\mathcal{N}_{\rho}(\partial\Omega))$ by the Calderón-Zygmund theory on VMO coefficients; see e.g. [6].
Then uvu-v is $\mathcal{A}$-harmonic in $\Omega$ and, since $u\in W^{2,\,2}(\Omega)$ is assumed,
@@
is the trace of a $W^{2,\,2}$-function on the boundary $\partial\Omega$ of a Lipschitz domain. Thus we have
@@
Now according to Lemma 3.2, by taking $\epsilon_{0}>0$ small enough so that $0<\omega_{0}\leq\epsilon_{0}\ll 1$, we can apply [11, Theorem 2.10] to vuv-u, together with fractional Sobolev embeddings, to conclude the proposition. ∎
@@
As $\Omega_{m}$ is $C^{2}$, the Hopf boundary lemma implies

### v4

@@ 1.2. Wulff potential and nontangential limits
One can generalize our approach for Theorem 1.1 to establish an anisotropic variant of Serrin’s overdetermined theorem for Lipschitz domains, with additional assumptions on $u$.
@@
Let $\Omega$ be a Lipschitz domain with a local Lipschitz constant $L>0$, $K$ be a (bounded) uniformly convex set of $C^{3}$ class with the principle curvatures of $\partial K$ bounded from above by $\kappa$ and from below by $\kappa^{-1}>0$, and $H$ be the associated Wulff potential. Assume that $u\in W^{1,\,2}(\mathbb{R}^{n})$ satisfies the anisotropic overdetermined system (1.4). Moreover, we suppose that $D^{2}u\in L^{n}$ and $|Du|\geq c_{0}>0$ in some neighborhood of $\partial\Omega$. Then there exists a constant $\epsilon_{0}=\epsilon_{0}(n,\,\kappa,\,\|K\|_{C^{3}})>0$ satisfying that, whenever $L\leq\epsilon_{0},$ one has
@@ 1.3. Alexandrov theorem and Maggi’s conjecture
for every $x\in\Omega$ and any $e\in\mathbb{S}^{n-1}$. Towards this, we first note that (1.3) implies
@@
As $\Omega_{m}$ is $C^{2}$, this implies
@@ 3. Preliminaries for the anisotropic case
$d\mu_{\mathcal{A}}(x)=\chi_{\Omega}(x)\sup\{d(z)|D\mathcal{A}(z)|^{2}\colon z\in B_{d(x)/4}(x)\}d\mathcal{L}(x),$
where $d(z)={\mathop{\mathrm{\,dist\,}}}(z,\,\partial\Omega)$, is a $C$-Carleson measure with respect to $\mathscr{H}^{n-1}|_{\partial\Omega}$, i.e., for any ball $B_{r}(y)$ centered at $y\in\partial\Omega$,
@@
$\mu_{\mathcal{A}}(B_{r}(y)\cap\Omega)\leq C\|D^{2}u\|^{2}_{L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)}r^{n-1},\quad 0<r<\rho$
where, for some $\rho>0$,
$\mathcal{N}_{\rho}(\partial\Omega):=\{z\in\mathbb{R}^{n}\colon{\mathop{\mathrm{\,dist\,}}}(z,\,\partial\Omega)<\rho\},\quad\omega_{0}=\|D^{2}u\|^{2}_{L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)},$
and
$0<C=C(n,\,c_{0},\,L,\,\kappa,\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})}).$
@@
Let $y\in\partial\Omega$ and $x\in B_{r}(y)$. By differentiating the equation in (1.4) in the $e_{k}$-direction, $1\leq k\leq n$, and letting $v=\partial_{k}u$, we obtain that $v$ satisfies, in the sense of distribution,
@@
where $\kappa$ is the curvature bound of $\partial K$ in Theorem 1.2. Then by summing over all $k$’s, as $|Du|\neq 0$ in a neighborhood of $\partial\Omega,$ it follows from the Schauder estimates in divergence form together with Poincaré inequality that, for any $0<t\leq\frac{d(x)}{4}$ with $B_{t}(x)$ disjoint from $\Omega_{0}\subset\subset\Omega$,
[…]
@@
and the (1)(-1)-homogeneity of $D^{3}V$ further gives that, for any $z\in B_{{d(x)}/4}(x)$,
@@
with uniformly bounded overlaps up to $C(n)$, covering $B_{r}(y)\cap\Omega$ so that, via applying (3.3) to $u$ on each $B_{d(x_{i})/4}(x_{i})$,
@@
$\leq$ $\ C(n,\,c_{0},\,\alpha,\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})})\sum_{i}d(x_{i})^{n-1}\left(\int_{B_{d(x_{i})/2}(x_{i})}|D^{2}u|^{n}\,dx\right)^{\frac{2}{n}}.$ (3.4)
Then when $n=2$, via $d(x_{i})\leq r$ we directly get
$\mu_{\mathcal{A}}({B_{r}(y)\cap\Omega})\leq C(c_{0},\,\alpha,\,\|D^{2}V\|_{C^{\alpha}(\mathbb{S}^{1})},\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{1})})\|D^{2}u\|^{2}_{L^{2}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)}r$
as desired.
When $n\geq 3$, we continue from (3.4) via Hölder’s inequality that, for $\beta=\frac{n(n-1)}{n-2},$
$\mu_{\mathcal{A}}({B_{r}(y)\cap\Omega})$
$\leq$ $\ C(n,\,c_{0},\,L,\,\alpha,\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})})\left(\sum_{i}\int_{B_{d(x_{i})/2}(x_{i})}|D^{2}u|^{n}\,dx\right)^{\frac{2}{n}}\left(\sum_{i}d(x_{i})^{\beta}\right)^{1-\frac{2}{n}}$
$\leq$ $\ C(n,\,c_{0},\,L,\,\alpha,\,\|D^{3}V\|_{L^{\infty}(\mathbb{S}^{n-1})})\|D^{2}u\|^{2}_{L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)}\left(\sum_{i}d(x_{i})^{\beta}\right)^{1-\frac{2}{n}}.$
Since there are at most $C(n,\,L)2^{l(n-1)}$-many balls in the collection with radius between $2^{-l}r$ and $2^{-l+1}r$ for each $l\in\mathbb{N}$, then
$\left(\sum_{i}d(x_{i})^{\beta}\right)^{1-\frac{2}{n}}\leq\left(\sum_{l}2^{l(n-1)}2^{-l(n-1)\frac{n}{n-2}}\right)^{1-\frac{2}{n}}r^{n-1}\leq C(n)r^{n-1},$
where $\frac{n}{n-2}+1<0$ was used. Thus, we obtain the desired estimate for $y\in\partial\Omega.$
@@
Note that the assumption of Theorem 1.2 yields the existence of $\rho>0$ for which
$D^{2}u\in L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)\quad\text{ and }\quad|Du|(z)\geq c_{0}>0\ \text{ for $z\in\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega$},$
@@
$\mathcal{N}_{\rho}(\partial\Omega):=\{z\in\mathbb{R}^{n}\colon{\mathop{\mathrm{\,dist\,}}}(z,\,\partial\Omega)<\rho\}.$
Then $Du\in VMO(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)$ directly by Sobolev embedding theorem.
@@
$\mathcal{A}(x)=D^{2}V(Du)\in VMO(\mathcal{N}_{\delta}(\partial\Omega)\cap\Omega),\quad D\mathcal{A}(z)=D^{3}V(Du(z))D^{2}u(z)\in L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)$
to $\mathbb{R}^{n}$ with $D\mathcal{A}(z)\in L^{n}(\mathcal{N}_{\rho}(\partial\Omega))$ and satisfying the ellipticity condition, and still denote the extended one by $\mathcal{A}$.
@@
Then since $D^{2}u\in L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega),$
${\rm div}({\mathcal{A}}(x)Dv)=\mathcal{A}D^{2}v+\partial_{i}\mathcal{A}_{ij}\partial_{j}v$
is uniformly elliptic, and $v\in W^{2,\,2}(\mathcal{N}_{\rho}(\partial\Omega))$ by the Calderón-Zygmund theory on VMO-coefficients; see e.g. [6].
Thus, uvu-v is $\mathcal{A}$-harmonic in $\Omega$ and, since $u\in W^{2,\,2}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)$,
@@
is the trace of a $W^{2,\,2}$-function on the boundary $\partial\Omega$ of a Lipschitz domain. Thus employing fractional Sobolev embeddings, we have
@@
Now according to Lemma 3.2, by taking $\rho>0$ small enough so that
$\|D^{2}u\|^{2}_{L^{n}(\mathcal{N}_{\rho}(\partial\Omega)\cap\Omega)}\leq\epsilon_{0}\ll 1$
we can apply [11, Theorem 2.10] to vuv-u to conclude the proposition. ∎
@@
As $\Omega_{m}$ is $C^{2}$, we get
@@
Conflict of interest statement: All authors declare that they have no conflicts of interest to disclose.